Lavage Rapide is a Canadian company that owns and operates a large automatic car wash facility near Montreal. The following table provides data concerning the company’s costs:

Fixed Cost
per Month
Cost per
Car Washed
Cleaning supplies $ 0.80
Electricity $ 1,100 $ 0.08
Maintenance $ 0.25
Wages and salaries $ 4,300 $ 0.20
Depreciation $ 8,400
Rent $ 1,900
Administrative expenses $ 1,600 $ 0.04

For example, electricity costs are $1,100 per month plus $0.08 per car washed. The company expects to wash 8,500 cars in August and to collect an average of $6.10 per car washed.


Required:

Prepare the company’s planning budget for August.

Revenue

Expenses:

Cleaning Supplies

Electricity

Maintanence

Wages and Salaries

Depreciation

Rent

Adminsitrative expenses

Total expense

Net operating income
Answer #1

ANSWER

Planning Budget
Budgeted Cars washed 8,500
Revenue 51,850 =8,500*6.10
Expenses:
Cleaning supplies 6,800 =8,500 *0.8
Electricity 1,780 =1,100 +(8,500 *0.08)
Maintenance 2,125 =8,500 *0.25
Wages and salaries 6,000 =4,300+(8,500 *0.20)
Depreciation 8,400
Rent 1,900
Administrative expenses 1,940 =1600+(8,500*0.04)
Total expenses 28,945
Net operating income 22,905

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************